MULTIVARIATE DISTRIBUTIONS IN STATISTICS

06–13.05.2003, Tampere

1. Introduction. Bivariate Normal Distribution.

Predominance of a normal distribution in statistics. Historical remarks. Some distribution properties. Regression and correlation. Characterizations. Simulation. Sampling distributions.

2. Multivariate Normal Distribution. Definitions. Basic distribution properties. Regression and correlation. Characterizations. Simulation. Sampling distributions.

3. Other Multivariate Distributions.

Literature

1. Tong, Y.L. The Multivariate Normal Distribution, Springer-Verlag, New York, 1990.

2. Johnson, N.L. and Kotz, S. Distributions in Statistics: Continuous Multivariate Distributions, John Wiley, New York, 1972.

3. Kotz, S., Balakrishnan, N. and Johnson, N.L. Continuous Multivariate Distributions. Vol. 1. Models and Applications, John Wiley, New York, 2000.

4. Morrison, D.F. Multivariate Statistical Methods, McGraw-Hill Kogakusha, 1976.

5. Pitman, J. Probability, Springer-Verlag, New York, 1992.

6. Jones, M.S. Multivariate t and beta distributions associated with the multivariate F distribution. Metrika, 2001, 54, 215-231.

7. Bratley, P., Fox, B.L. and Schrage, L.E. A Guide to Simulation, Springer-Verlag, New York, 1987.

8. Fishman, G.S. Monte-Carlo. Concepts, Algorithms, and Applications, Springer-Verlag, New York, 1996. 9 . Gentle, J.E. Random Number Generation and Monte-Carlo Methods, Springer-Verlag, New York, 1998.